OpenStudy (anonymous):

After (x+1)/(3)-(x+2)/(2x) have been combined using the Least Common Denominator of 6x, what is the numerator?

OpenStudy (mertsj):

\[\frac{x+1}{3}-\frac{x+2}{2x}=\frac{2x(x+1)}{6x}-\frac{3(x+2)}{6x}=\frac{2x^2-x-6}{6x}\]

OpenStudy (anonymous):

thank you!

OpenStudy (mertsj):

yw